Circular No. 6736

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOTICE TO CONTRIBUTORS Over the course of twelve hours on Sept. 6 UT the Central Bureau

received two reports—purportedly independent—of the CCD discovery of an alleged supernova in NGC 772. Each indicated that the object was of mag 15–16, but the first was offset ∼ 2′ northeast and the second ∼ 2′ northwest of the center of the galaxy; the first observer did not record the observation time, but this image was evidently obtained ∼ 26 hr before the second observer’s image was obtained. Each observer noted that the alleged supernova was not present on images from earlier nights and that it did not move over the course of an hour or two; furthermore, the first observer remarked that two different checks for known minor planets in the vicinity did not reveal a match. The second observer indicated that this report had also been sent to selected other observers, with a note that “spectroscopic confirmation was requested”. On receipt of the information from the first observer, a check at the Central Bureau immediately revealed that, although the observation time had not been specified, the alleged supernova northeast of the galaxy was probably minor planet (5240), then very close to its morning stationary point and in fact still moving very slowly eastward. This information was relayed to the observer. Nevertheless, within six hours of this (i.e., still two hours before the second observer’s image was obtained), a copy of the first observer’s image was present in the World Wide Web, together with the mention (though without details) of an apparent confirmation by the webmaster. Measurement of the WWW image at the Central Bureau clearly confirmed the indicated identity of the “supernova” with (5240) and allowed the observation time to be specified. In the Bureau’s message to the first observer, it was also mentioned that minor planet (1887), comparable to (5240) in brightness, was near the galaxy. Examination of the WWW image showed that (1887) would have been just off the southern edge at the time of observation. Nevertheless, (1887) was moving northward at five times the rate of (5240), and it had already begun its westward motion. On receipt of the second report of an alleged supernova, with the indication that it was 125′′ west and 75′′ north of the center of the galaxy, it was then immediately clear that this observation in fact referred to (1887).

The above account is given in some detail because it allows several points to be made. These will be discussed on IAUC 6737, together with a proposed remedy for the problem.

1997 September 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden